### v6

@@ Abstract
Bilevel optimization reveals the inner structure of otherwise oblique optimization problems, such as hyperparameter tuning, neural architecture search, and meta-learning. A common goal in bilevel optimization is to minimize a hyper-objective that implicitly depends on the solution set of the lower-level function. Although this hyper-objective approach is widely used, its theoretical properties have not been thoroughly investigated in cases where the lower-level functions lack strong convexity. In this work, we first provide hardness results to show that the goal of finding stationary points of the hyper-objective for nonconvex-convex bilevel optimization can be intractable for zero-respecting algorithms. Then we study a class of tractable nonconvex-nonconvex bilevel problems when the lower-level function satisfies the Polyak-Łojasiewicz (PL) condition. We show a simple first-order algorithm can achieve complexity bounds of $\tilde{\mathcal{O}}(\epsilon^{-2})$, $\tilde{\mathcal{O}}(\epsilon^{-4})$ and $\tilde{\mathcal{O}}(\epsilon^{-6})$ in the deterministic, partially stochastic, and fully stochastic setting respectively.
@@ 1 Introduction
The PL condition can pose nontrivial challenges since we have neither Equation (3) nor (4) in this case. Researchers have provided novel analyses in this case. Xiao et al. (2023) proposed a Hessian-vector-product-based method GALET with non-asymptotic convergence to KKT points of the gradient-based reformulated problem (Equation (11)) when $g$ is PL in $y$. Kwon et al. (2024b) proved the differentiability of $\varphi(x)$ when the penalty function $\sigma f+g$ is uniformly PL in $y$ for all $\sigma$ in the neighborhood of zero, and showed a proximal variant of F2BA, which we call Prox-F2BA, can find a stationary point of $\varphi(x)$. Based on their differentiability result, we show that GALET also converges to a stationary point of $\varphi(x)$ in Appendix B.
Although Prox-F2BA (Kwon et al., 2024b) has been shown to converge to an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls, the rate is worse than the ${\mathcal{O}}(\epsilon^{-2})$ rate of gradient descent on nonconvex single-level optimization. Therefore, listed as an important future direction, Kwon et al. (2024b) asked the question below.
Can one achieve the ${\mathcal{O}}(\epsilon^{-2})$ rate for nonconvex-PL bilevel problems with gradient oracles?
We give a positive answer to this question. Interestingly, we prove that the original F2BA (Chen et al., 2023) can already achieve this fast convergence rate. Our improvement over Kwon et al. (2024b) comes from establishing a tighter bound on the smoothness constant of $\varphi(x)$.
@@
1st Prox-F2BA (b) $\tilde{\mathcal{O}}(\kappa^{p_{1}}\epsilon^{-3})$ $\tilde{\mathcal{O}}(\kappa^{p_{2}}\epsilon^{-5})$ $\tilde{\mathcal{O}}(\kappa^{p_{3}}\epsilon^{-7})$ Kwon et al. (2024b)
1st F2BA $\tilde{\mathcal{O}}(\kappa^{4}\epsilon^{-2})$ $\tilde{\mathcal{O}}(\kappa^{5}\epsilon^{-4})$ $\tilde{\mathcal{O}}(\kappa^{11}\epsilon^{-6})$ This Paper
  • (a) Although Xiao et al. (2023) did not provide the dependency on $\kappa$ in the complexity, we can calculate by the way in our Remark B.1. Their analysis additionally requires the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero, but such assumption seems to be redundant by our Remark B.2.
  • (b) We use $p_{1},p_{2},p_{3}$ to denote the polynomial dependency in $\kappa$ since they are not provided by Kwon et al. (2024b).
@@
Under regular conditions as in Kwon et al. (2024b), F2BA can provably find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-2})$ first-order oracle calls for nonconvex-PL bilevel problems.
We also extend our analysis to stochastic bilevel problems when an algorithm only has access to a noisy estimator of $\nabla f$ and $\nabla g$. Under the partially stochastic setting when the noise is only in $\nabla g$, we prove the stochastic F2BA has the $\tilde{\mathcal{O}}(\epsilon^{-4})$ first-order oracle complexity, matching SGD for single-level nonconvex minimization (Arjevani et al., 2023). Under the more general fully stochastic setting when noise appears both in $\nabla f$ and $\nabla g$, we prove the method has an $\tilde{\mathcal{O}}(\epsilon^{-6})$ complexity.
Compared with Xiao et al. (2023), our deterministic F2BA achieves the same $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate without the assistance of Hessian-vector-product oracles. Our method also has a better dependency on $\kappa$ since we do not use a squared trick on the Hessian of $g$. And we additionally study stochastic problems that have not been studied by Xiao et al. (2023). Compared with Kwon et al. (2024b), we strictly improve the complexities for both deterministic and stochastic cases under the same assumptions. We compare our results with prior works in Table 1, and leave a more detailed introduction of related works in Appendix A.
@@ 3.2 Stationary Points May be Intractable to Find
The above zero-respecting algorithm class for bilevel problems subsumes many known algorithms when applied to Problem (5), including: AID (Ghadimi and Wang, 2018), ITD (Ji et al., 2021), GALET (Xiao et al., 2023), (Prox)-F2BA (Kwon et al., 2024b, 2023; Chen et al., 2023), FdeHBO (Yang et al., 2024), BGS-Opt (Arbel and Mairal, 2022), BDA (Liu et al., 2020), BVFIM (Liu et al., 2021a), PDBO (Sow et al., 2022), SLM (Lu, 2023), LV-HBA (Yao et al., 2024).
@@ 4 Positive Results for Lower-Level Functions Satisfying PL Conditions
As we have shown finding small hyper-gradients of nonconvex-convex bilevel problems is intractable for nonconvex-convex bilevel problems, we turn our attention to the tractable cases. Motivated by recent works (Shen and Chen, 2023; Kwon et al., 2024b; Arbel and Mairal, 2022), we study the case when the lower-level problem satisfies the PL condition.
@@ 4.1 The Assumptions for Nonconvex-PL Bilevel Problems
Without the typical lower-level strong convexity assumption, it is difficult to directly analyze the implicit gradient as Equation (4) since we can not directly use the implicit function theorem. Recently, Kwon et al. (2024b) proposed a novel way to study the differentiability of $\varphi(x)$ for nonconvex-PL bilevel problems. Instead of directly studying the original hyper-objective $\varphi(x)$, they studied the following value-function penalized hyper-objective as a bridge:
@@
To guarantee the differentiability of $\varphi_{\sigma}(x)$, Kwon et al. (2024b) assumed the PL condition not only holds for $g$, but also holds for the penalty function $h_{\sigma}=\sigma f+g$ uniformly for all $\sigma$ near zero. In addition to this assumption, the authors also imposed other standard smoothness assumptions which are typically required in previous works. These assumptions, as stated in Kwon et al. (2024b), are formally presented below.
@@
    The upper-level function $f(x,y)$ has $\rho_{f}$-Lipschitz Hessians;
@@
Note that all the assumptions are the same as Kwon et al. (2024b), except (a) may seem different from the Prox-EB assumption (Assumption 1 in Kwon et al. (2024b)). However, in Appendix C we show these two assumptions can imply each other. Our narrative uses the PL condition and is therefore more convenient. As a by-product, we show the proximal operator in Kwon et al. (2024b) is unnecessary, and the original F2BA (Chen et al., 2023) can also converge under PL conditions.
Because we focus on the behavior when $\sigma$ is close to zero, whenever we mention $\sigma$ in the context, we always assume that $\sigma\in[0,\overline{\sigma}]$ even when this condition is not explicitly stated.
@@ 4.2 F2BA Can also be Applied to Nonconvex-PL Bilevel Problems
$\nabla\varphi_{\sigma}(x)=\nabla_{x}f(x,y)+\frac{\nabla_{x}g(x,y_{\sigma}^{*}(x))-\nabla_{x}g(x,y^{*}(x))}{\sigma},$ (7)
@@
Secondly, the stability of $Y_{\sigma}^{*}$ under perturbations of $\sigma$ by Lemma 2 implies the stability of $\varphi_{\sigma}$ and $\nabla\varphi_{\sigma}$ by invoking the result by Kwon et al. (2024b).
Lemma 4 (Kwon et al. (2024b)).
@@
These two lemmas make it reasonable to apply the F2BA (Algorithm 2), which applies gradient descent on $\varphi_{\sigma}(x)$ according to Equation (7). Lemma 4 shows that to find an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi(x)$, it suffices to finds an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi_{\sigma}(x)$ for $\sigma={\mathcal{O}}(\epsilon)$. Note that gradient descent can find an $\epsilon$-stationary point of a nonconvex $L$-smooth function within ${\mathcal{O}}(L\epsilon^{-2})$ complexity and that $\varphi_{\sigma}(x)$ is ${\mathcal{O}}(\sigma^{-1})$-smooth in the worst-case. Such an analysis illustrates an $\tilde{\mathcal{O}}(\sigma^{-1}\epsilon^{-2})=\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity as Kwon et al. (2024b).
@@
4.3 Achieving the $\tilde{\mathcal{O}}(\epsilon^{-2})$ Rate for the Deterministic Case
@@
The above complexity matches gradient descent for single-level nonconvex minimization by Carmon et al. (2021), except we require $f$ to be $\rho_{f}$-Hessian Lipschitz in Assumption 4.1d. However, our dependency on $\rho_{f}$ is only logarithmic, rather than polynomial.
@@
Here we note some small differences between the assumptions in Theorem 6 and those of the constructed hard instance in Theorem 2.
One difference is that the hard instance only satisfies the Lipschitz-type conditions in ${\mathcal{X}}\times{\mathbb{R}}^{d_{y}}$, where ${\mathcal{X}}$ is the sublevel set of $x_{0}$. However, indeed, the upper bound also only requires the Lipschitz-type conditions in the sublevel set by the descent lemma $\varphi(x_{t+1})\leq\varphi(x_{t})$.
Another difference is that the upper bound requires $f$ to have $\rho_{f}$-Lipschitz Hessians while the Hessian Lipschitz constant in the hard instance may depend on $d_{y}$. However, the upper bound only has a logarithmic dependency on $\rho_{f}$. Such a difference is also mild.
Besides the above two differences, the only remaining difference is that the lower-level function in the hard instance is only convex, while Assumption 4.1a required by the upper bound subsumes local strong convexity. Such a difference leads to a separation of intractability and tractability.
@@ 4.4 The Extensions of F2BA to the Stochastic Case
$\begin{split}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\eta\asymp\ell^{-1}\kappa^{-3},\quad\sigma\asymp\min\left\{\frac{R}{\kappa},~{}\frac{\epsilon}{\ell\kappa^{3}},~{}\frac{L_{g}}{L_{f}},~{}\frac{\rho_{g}}{\rho_{f}},~{}\overline{\sigma}\right\},\\ &\tau=\frac{1}{\sigma L_{f}+L_{g}},\quad{B\asymp\frac{L_{g}\left(\sigma^{2}M_{f}^{2}+M_{g}^{2}\right)}{\mu\sigma^{2}\epsilon^{2}}},\quad K_{t}\asymp\frac{L_{g}}{\mu}\log\left(\frac{L_{g}^{3}\delta_{t}}{\mu\sigma^{2}\epsilon^{2}}\right),\end{split}$ (8)
@@
$\begin{cases}{\mathcal{O}}(\ell\kappa^{4}\epsilon^{-2}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}})),&M_{f}=0,M_{g}=0;\\ {\mathcal{O}}(\ell\kappa^{5}\epsilon^{-4}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}})),&M_{f}>0,M_{g}=0;\\ {{\mathcal{O}}(\ell^{3}\kappa^{11}\epsilon^{-6}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}}))},&M_{f}>0,M_{g}>0.\end{cases}$
In the deterministic case $(M_{f}=0,M_{g}=0)$, this result recovers the $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate by Algorithm 2. In the partially stochastic case ($M_{f}>0,M_{g}=0$), the $\tilde{\mathcal{O}}(\epsilon^{-4})$ complexity also matches SGD (Arjevani et al., 2023). In the fully stochastic case ($M_{f}>0,M_{g}>0$), our $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper bound is also better than the $\tilde{\mathcal{O}}(\epsilon^{-7})$ upper bound by Kwon et al. (2024b).

### v8

@@ Abstract
Bilevel optimization reveals the inner structure of otherwise oblique optimization problems, such as hyperparameter tuning, neural architecture search, and meta-learning. A common goal in bilevel optimization is to minimize a hyper-objective that implicitly depends on the solution set of the lower-level function. Although this hyper-objective approach is widely used, its theoretical properties have not been thoroughly investigated in cases where the lower-level functions lack strong convexity. In this work, we first provide hardness results to show that the goal of finding stationary points of the hyper-objective for nonconvex-convex bilevel optimization can be intractable for zero-respecting algorithms. Then we study a class of tractable nonconvex-nonconvex bilevel problems when the lower-level function satisfies the Polyak-Łojasiewicz (PL) condition. We show a simple first-order algorithm can achieve complexity bounds of $\tilde{\mathcal{O}}(\epsilon^{-2})$, $\tilde{\mathcal{O}}(\epsilon^{-4})$ and $\tilde{\mathcal{O}}(\epsilon^{-6})$ in the deterministic, partially stochastic, and fully stochastic setting respectively. The complexities in the first two cases are optimal up to logarithmic factors.
@@ 1 Introduction
The PL condition can pose nontrivial challenges since we have neither Equation (3) nor (4) in this case. Researchers have provided novel analyses in this case. Xiao et al. (2023) proposed a Hessian-vector-product-based method GALET with non-asymptotic convergence to KKT points of the gradient-based reformulated problem (Equation (11)) when $g$ is PL in $y$. Kwon et al. (2024) proved the differentiability of $\varphi(x)$ when the penalty function $\sigma f+g$ is uniformly PL in $y$ for all $\sigma$ in the neighborhood of zero, and showed a proximal variant of F2BA, which we call Prox-F2BA, can find a stationary point of $\varphi(x)$. Based on their differentiability result, we show that GALET also converges to a stationary point of $\varphi(x)$ in Appendix B.
Although Prox-F2BA (Kwon et al., 2024) has been shown to converge to an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls, the rate is worse than the ${\mathcal{O}}(\epsilon^{-2})$ optimal rate of gradient descent on nonconvex single-level optimization. Therefore, listed as an important future direction, Kwon et al. (2024) asked the question below.
Can one achieve the (near)-optimal rate for nonconvex-PL bilevel problems with gradient oracles?
We give a positive answer to this question. We show the F2BA (Chen et al., 2023) can already achieve this fast convergence rate with a sharp analysis. Our improvement over Kwon et al. (2024) comes from establishing a tighter bound on the smoothness constant of $\varphi(x)$.
@@
1st Prox-F2BA (b) $\tilde{\mathcal{O}}(\kappa^{p_{1}}\epsilon^{-3})$ $\tilde{\mathcal{O}}(\kappa^{p_{2}}\epsilon^{-5})$ $\tilde{\mathcal{O}}(\kappa^{p_{3}}\epsilon^{-7})$ Kwon et al. (2024)
1st F2BA $\tilde{\mathcal{O}}(\kappa^{4}\epsilon^{-2})$ $\tilde{\mathcal{O}}(\kappa^{6}\epsilon^{-4})$ $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$ This Paper
  • (a) Although Xiao et al. (2023) did not provide the dependency on $\kappa$ in the complexity, we can calculate by the way in our Remark B.1. Their analysis additionally requires the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero, which makes $\nabla_{yy}^{2}g(x,y)$ have a constant rank.
  • (b) We use $p_{1},p_{2},p_{3}$ to denote the polynomial dependency in $\kappa$ since they are not provided by Kwon et al. (2024).
@@
Under regular conditions as Kwon et al. (2024), F2BA can provably find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-2})$ first-order oracle calls for nonconvex-PL bilevel problems.
We also extend our analysis to stochastic bilevel problems when an algorithm only has access to a noisy estimator of $\nabla f$ and $\nabla g$. Under the partially stochastic setting when the noise is only in $\nabla g$, we prove the stochastic F2BA has the $\tilde{\mathcal{O}}(\epsilon^{-4})$ first-order oracle complexity, which is also near-optimal for stochastic optimization (Arjevani et al., 2023). Under the more general fully stochastic setting when noise appears both in $\nabla f$ and $\nabla g$, we prove the method has an $\tilde{\mathcal{O}}(\epsilon^{-6})$ complexity.
Compared with Xiao et al. (2023), our deterministic F2BA achieves the same $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate without the assistance of Hessian-vector-product oracles. Our method also has a better dependency on $\kappa$ since we do not use a squared trick on the Hessian of $g$. And we additionally study stochastic problems that have not been studied by Xiao et al. (2023). Compared with Kwon et al. (2024), we strictly improve the complexities for both deterministic and stochastic cases under the same assumptions. We compare our results with prior works in Table 1, and leave a more detailed introduction of related works in Appendix A.
@@ 3.2 Stationary Points May be Intractable to Find
The above zero-respecting algorithm class for bilevel problems subsumes many known algorithms when applied to Problem (5), including: AID (Ghadimi and Wang, 2018), ITD (Ji et al., 2021), GALET (Xiao et al., 2023), (Prox)-F2BA (Kwon et al., 2024, 2023; Chen et al., 2023), FdeHBO (Yang et al., 2024), BGS-Opt (Arbel and Mairal, 2022), BDA (Liu et al., 2020), BVFIM (Liu et al., 2021a), PDBO (Sow et al., 2022), SLM (Lu, 2023), LV-HBA (Yao et al., 2024).
@@ 4 Positive Results for Lower-Level Functions Satisfying PL Conditions
As we have shown finding small hyper-gradients of nonconvex-convex bilevel problems is intractable for nonconvex-convex bilevel problems, we turn our attention to the tractable cases. Motivated by recent works (Shen and Chen, 2023; Kwon et al., 2024; Arbel and Mairal, 2022), we study the case when the lower-level problem satisfies the PL condition.
@@ 4.1 The Assumptions for Nonconvex-PL Bilevel Problems
Without the typical lower-level strong convexity assumption, it is difficult to directly analyze the implicit gradient as Equation (4) since we can not directly use the implicit function theorem. Recently, Kwon et al. (2024) proposed a novel way to study the differentiability of $\varphi(x)$ for nonconvex-PL bilevel problems. Instead of directly studying the original hyper-objective $\varphi(x)$, they studied the following value-function penalized hyper-objective as a bridge:
@@
To guarantee the differentiability of $\varphi_{\sigma}(x)$, Kwon et al. (2024) assumed the PL condition not only holds for $g$, but also holds for the penalty function $h_{\sigma}=\sigma f+g$ uniformly for all $\sigma$ near zero. In addition to this assumption, the authors also imposed other standard smoothness assumptions which are typically required in previous works. These assumptions, as stated in Kwon et al. (2024), are formally presented below.
@@
    The upper-level function $f(x,y)$ has $\rho_{f}$-Lipschitz Hessians in $y$, i.e. $\nabla_{xy}^{2}f$ and $\nabla_{yy}^{2}f$ are $\rho_{f}$-Lipschitz continuous;
@@
Because we focus on the behavior when $\sigma$ is close to zero, whenever we mention $\sigma$ in the context, we always assume that $\sigma\in[0,\overline{\sigma}]$ even when this condition is not explicitly stated. Note that all the assumptions are the same as Kwon et al. (2024), except Assumption 4.1(a) may seem different from the Prox-EB assumption (Assumption 1 in Kwon et al. (2024)). In Appendix C we show these two assumptions can imply each other in the unconstrained case. Our narrative uses the PL condition and is therefore more convenient. As a by-product, we show the proximal operator in Kwon et al. (2024) is unnecessary, and the original F2BA (Chen et al., 2023) can also converge under PL conditions. In Assumption 4.1(d), Kwon et al. (2024) simplily assumes $f(x,y)$ has Lipschitz Hessians, but we note that their analysis only requires $\nabla_{xy}^{2}f$ and $\nabla_{yy}^{2}f$ are Lipschitz continuous. We use this refined assumption because the $\Omega(\epsilon^{-2})$ lower bound function for finding an $\epsilon$-stationary point of $f(x)$ does not have Lipschitz continuous Hessians in $x$, so we also do not assume $\nabla_{xx}^{2}f$ is Lipschitz continuous in our upper bounds.
Remark 4.2
Assumption 4.1(a) can be also replaced by $g(x,y)$ satisfies $\mu$-PL condition with a unique minimizer $y^{*}(x)$ and $\nabla_{yy}^{2}g(x,y^{*}(x))$ is non-singular as Huang (2023, 2024). The analysis would be almost the same, but it is an easier case as discussed in Section 3.2.1 (Kwon et al., 2024). In this paper, we focus on the more challenging case as in (Kwon et al., 2024).
@@ 4.2 F2BA Can also be Applied to Nonconvex-PL Bilevel Problems
$\nabla\varphi_{\sigma}(x)=\nabla_{x}f(x,y_{\sigma}^{*}(x))+\frac{\nabla_{x}g(x,y_{\sigma}^{*}(x))-\nabla_{x}g(x,y^{*}(x))}{\sigma},$ (7)
@@
Secondly, the stability of $Y_{\sigma}^{*}$ under perturbations of $\sigma$ by Lemma 2 implies the stability of $\varphi_{\sigma}$ and $\nabla\varphi_{\sigma}$ by invoking the result by Kwon et al. (2024).
Lemma 4 (Kwon et al. (2024)).
@@
These two lemmas make it reasonable to apply the F2BA (Algorithm 2), which applies gradient descent on $\varphi_{\sigma}(x)$ according to Equation (7). Lemma 4 shows that to find an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi(x)$, it suffices to finds an ${\mathcal{O}}(\epsilon)$-stationary point of $\varphi_{\sigma}(x)$ for $\sigma={\mathcal{O}}(\epsilon)$. Note that gradient descent can find an $\epsilon$-stationary point of a nonconvex $L$-smooth function within ${\mathcal{O}}(L\epsilon^{-2})$ complexity and that $\varphi_{\sigma}(x)$ is ${\mathcal{O}}(\sigma^{-1})$-smooth in the worst-case. Such an analysis illustrates an $\tilde{\mathcal{O}}(\sigma^{-1}\epsilon^{-2})=\tilde{\mathcal{O}}(\epsilon^{-3})$ complexity as Kwon et al. (2024).
@@
4.3 Achieving the Near-Optimal Rate for the Deterministic Case
@@
The above complexity matches the optimal rate gradient descent for single-level nonconvex minimization by Carmon et al. (2021), except for an additional logarithmic factor.
@@
One small difference between the assumptions in Theorem 6 and those of the constructed hard instance in Theorem 2. is that the hard instance only satisfies the Lipschitz-type conditions in ${\mathcal{X}}\times{\mathbb{R}}^{d_{y}}$, where ${\mathcal{X}}$ is the sublevel set of $x_{0}$. However, indeed, the upper bound also only requires the Lipschitz-type conditions in the sublevel set by the descent lemma $\varphi(x_{t+1})\leq\varphi(x_{t})$.
@@ 4.4 The Extensions of F2BA to the Stochastic Case
$\begin{split}&~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\eta\asymp\ell^{-1}\kappa^{-3},\quad\sigma\asymp\min\left\{\frac{R}{\kappa},~{}\frac{\epsilon}{\ell\kappa^{3}},~{}\frac{L_{g}}{L_{f}},~{}\frac{\rho_{g}}{\rho_{f}},~{}\overline{\sigma}\right\},\\ &\tau=\frac{1}{\sigma L_{f}+L_{g}},\quad{B\asymp\frac{L_{g}^{2}\left(\sigma^{2}M_{f}^{2}+M_{g}^{2}\right)}{\mu^{2}\sigma^{2}\epsilon^{2}}},\quad K_{t}\asymp\frac{L_{g}}{\mu}\log\left(\frac{L_{g}^{3}\delta_{t}}{\mu\sigma^{2}\epsilon^{2}}\right),\end{split}$ (8)
@@
$\tilde{\mathcal{O}}(\kappa TB)=\begin{cases}{\mathcal{O}}(\ell\kappa^{4}\epsilon^{-2}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}})),&M_{f}=0,M_{g}=0;\\ {\mathcal{O}}(\ell\kappa^{6}\epsilon^{-4}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}})),&M_{f}>0,M_{g}=0;\\ {{\mathcal{O}}(\ell^{3}\kappa^{12}\epsilon^{-6}\log(\nicefrac{{\rho_{f}\ell\kappa}}{{\epsilon}}))},&M_{f}>0,M_{g}>0.\end{cases}$
In the deterministic case $(M_{f}=0,M_{g}=0)$, this result recovers the $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate by Algorithm 2. In the partially stochastic case ($M_{f}>0,M_{g}=0$), the $\tilde{\mathcal{O}}(\epsilon^{-4})$ complexity is also near-optimal (Arjevani et al., 2023). In the fully stochastic case ($M_{f}>0,M_{g}>0$), our $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper bound is also better than the $\tilde{\mathcal{O}}(\epsilon^{-7})$ upper bound by Kwon et al. (2024).